### v2

# Richard A Berk

Abstract
This paper provides forecasts of rare and extreme heat waves coupled with estimates of uncertainty. Both rest in part on gaining a better understanding of the similarities and differences between hot days under normal circumstances and rare, extreme heat waves. We analyze AIRS data from the American Pacific Northwest and AIRS data from the Phoenix, Arizona region. Forecasting accuracy is excellent for the Pacific Northwest and is replicated for Phoenix. Uncertainty is estimated with nested conformal prediction sets, provably valid even for finite samples. The prediction sets are precise, and the coverage probabilities high. Complications from endogenous sampling are discussed.
Keywords: heat waves, forecasting, AIRS data, machine learning, genetic algorithms, endogenous sampling
1 Introduction
Extreme heat waves are relatively rare events. But of late, extreme heat waves seem to be increasing in frequency and intensity (Mann et al., 2018). Schmidt (2024) argues that “heat anomalies” may mean that climate science has underestimated the speed at which disastrous global warming will materialize. Many of the consequences for public health and local ecosystems are already troubling (Barriopedro et al., 2023). Some projected consequences are dire (Khatana et al., 2024). The mass media also has taken note. From the Economist, for example, “Deadly heat is increasingly the norm, not an exception to it.” (Economist, June 6, 2024).
Existing studies seeking to understand extreme heat waves have so far produced mixed results (Mann et al., 2018; McKinnon and Simpson, 2022, Fischer et al., 2023; Zhang et al., 2024; Li et al., 2024). Part of the problem is that during the warm-weather months, sequences of hot days have long been “normal.” These spells have some of the features of heat waves that some have linked to climate change. But they arguably are not the rare and extreme heat waves about which there has been so much recent concern.
In this manuscript, the primary goal is to accurately forecast rare and extreme heat waves. To that end, we use novel methods in an effort to help distinguish between usual summer days and the rare, extreme heat waves, perhaps induced by climate change. The data used are a Level 3 product of the NASA Atmospheric Infrared Sounder (AIRS) instrument aboard the Aqua spacecraft (Aumann et al., 2003; Pagano et al., 2010). Observations from the June 2021 heat wave in the North American, Pacific Northwest (PNW) provide an initial forecasting target. Observations from the 2023 July heat wave in the greater Phoenix, Arizona area allow for a possible replication. For both sites, comparisons constructed from “usual” summer days are taken from other months or other years for which no credible claims of extreme temperature episodes have been made. We achieve uncertainty guarantees using nested conformal prediction sets (Kuchibhotla and Berk, 2023).
Section 2 addresses the methodological challenges raised by efforts to forecast extreme heat waves, typically characterized by temperatures that likely fall in the far right tail of temperature distributions; they represent rare events that commonly cause serious data sparsity problems. Section 3 provides a grounded discussion of the remotely sensed, AIRS grid cell data to be used and in Section 4, the research designs that can work well with rare events like extreme heat waves are introduced. Substantial progress is made on data sparsity and on design-based adjustments to mitigate the impact of nuisance variables. Section 5 uses gain scores to examine the marginal, ground-level temperature distributions of an actual heat wave compared to a faux heat wave. The distributions are rather different. Section 6 builds on several variants of crossover designs, with grid cells as the study units, to determine how well machine learning can fit ground-level temperatures during an extreme heat wave compared to ground-level temperatures from ordinary hot summer days. We find that machine learning performance is good, and there is some evidence that the dominant predictors of heat wave temperatures differ somewhat from the dominant predictors of “normal” hot summer temperatures. Section 7 introduces a genetic algorithm to construct a synthetic population of ideal-type, extreme heat wave grid cells, and a synthetic population of ideal-type, ordinary high temperature summer grid cells. Comparisons between the predictor values in these different populations enable a simple form of feature engineering resulting in variable selection. Section 8 uses the selected predictors to forecast a binary response variable denoting whether or not a grid cell experienced an extreme heat wave. Forecasting skill is very good, and nested conformal prediction sets provide provably valid estimates of uncertainty. Section 9 reports a replication of the Pacific Northwest results using an extreme heat wave in the greater Phoenix area. As one would expect, the findings differ a bit, but the overall comparability of the broad conclusions is promising. Section 10 returns to the structure of our research designs in the context of endogenous sampling. We consider the degree to which our efforts to overcome the rare event problem have introduced difficulties in the reported results. We suggest that there is promise in a reformulation from minimizing loss to minimizing risk. Section 11 offers a brief summary of the results and some overall conclusions.
2 Methodological Challenges
In the pages ahead, we use the term “heat wave” to denote extreme heat waves that are relatively rare and arguably enabled by climate change. We use the term “heat dome” as well to capture some instructive climate science thinking (e.g., Jain at el., 2024). Within this framing, there can be many definitions that differ at least in important details (Perkins, 2015). The definitions we use are provisional, constrained in part by available information in the AIRS data.
Almost regardless of how heat waves are defined, a major statistical challenge is data sparsity because extreme heat waves are relatively unusual. One consequence is that very accurate but trivial forecasts can be guaranteed. If in most remotely sensed grid cells most of the time, there is no heat wave, a forecast of no heat wave will almost always be correct without needing to bother with predictors. When predictors are available, they typically will fail to improve the fit or forecasts because both are nearly perfect already. A more subtle consequence is that the design space defined by the predictors will be empty in many regions, making it nearly impossible to estimate well many kinds of nonlinear functions. The fallback of linear statistical methods can be overmatched and miss very important heat wave characteristics. Some consequences of these and related difficulties are addressed later.
Current global climate models (GCMs) do not appear to offer much help. For the spatial scales at which global climate models function, heat waves are generally below the spatial scales that can be well resolved. Various forms of downscaling seem necessary to construct regional climate models (RCMs), but these introduce a number of new challenges (Gebrechorkos et al., 2023; Lafferty and Sriver, 2023; Schneider et al., 2023). As Gettelman and Rood (2016) write, “For these [smaller scale] processes, we often must use statistical treatments to match observations to functions that can be used to describe the behavior. Some processes can be represented by basic laws, other processes must (at the scale of a climate model) be represented by statistical relationships that are only as good as our observations.” To this we add: and only as good as the statistical models themselves (Breiman, 2001b). Moreover, capitalizing on “basic laws” depends on knowing the relevant physics in particular, complex settings, getting boundary conditions from a GCM that are approximately correct, and having access to significant computational resources. And even then, the spatial scale may be too coarse. The analyses to follow suggest that certain quasi-experimental research designs coupled with that machine learning algorithms can help.
There are also difficulties in how best to represent uncertainty. Because there is apparently no consensus model for extreme heat waves, there is no widely acceptable way to formulate a generative model for how the observations materialize. Without such a formulation, the usual uncertainty calculations risk being little more than hand waving. Gettelman and Rood (2016) explain that in climate science, uncertainty is caused by (1) model uncertainty, (2) scenario uncertainty and (3) initial condition uncertainty. In statistics and econometrics, the first is called “model misspecification” (Freedman, 2009). The second is closely related to various forms of stationarity in statistical time series analysis (Box et al., 2016). The third overlaps in statistics with the literature on errors in variables (Fuller, 1987). How these and other statistical and econometric formulations map to those in climate science is largely unresolved.
Further confusion sometimes is introduced by different meanings of the term “model” (Berk et al., 2024). In the natural sciences, a model typically is intended to facilitate explanation and, often, causal inference. The focus is on the manner in which some empirical phenomenon works, and understanding is commonly the dominant goal. Such models necessarily introduce simplifications to further understanding. There also are various other ways models systematically can be right or wrong.
In addition, there can be confusion when an algorithm is called a “model.” “An algorithm is nothing more than a very precisely specified series of instructions for performing some concrete task” (Kearns and Roth, 2019). The goal is to perform that task well without necessarily representing the underlying subject-matter mechanisms. Algorithms are neither right nor wrong; as a technical matter, there is no such thing as a misspecified algorithm. Either an algorithm performs in a satisfactory fashion or it does not. For example, a pattern recognition algorithm might be used to identify malignancies in lung X-rays and often can perform better at that task than a radiologist (Li et al., 2022). But, there is in the algorithm no necessary explanatory or causal content for how the malignant tissue originated and grew.
Frequentist statisticians often use the term “model” in yet another way (Berk et al., 2024). A model usually conveys how the data were generated. Uncertainty commonly begins with random variables and, therefore, the means by which the data were realized. Sometimes the data are realized through design-based approaches, as in various forms of probability sampling used estimate parameters of populations or in random assignment used to estimate causal effects in experiments. Then, a climate science narrative usually is a subject-matter overlay. Sometimes the data are assumed to be realized through processes characterized by subject-matter models as just described. Such models often build on climate science theory. Climate science examples include applications of the generalized extreme value distribution and all manner of linear regression, sometimes hand tailored for time series. All of the usual concerns about model misspecification commonly arise (Berk et al., 2019). In the pages ahead, we will try to make clear when using models, which kind of model we have in mind.
A more subtle potential problem with all models is that as a formal matter, a model should be fully determined before the data are examined. Yet, much model specification is based on model foraging, often as recommended practice. Widely cited treatments of the generalized extreme value distribution (Coles, 2001; Gilleland and Katz, 2016), for instance, provide a variety of statistical tools for empirical determination of which special case of the umbrella formulation that should be employed: the Gumbel, Fréchet or Weibull distribution (Kotz et al., 2000). The result is sanctioned cherry picking, which can invalidate all statistical inference (Leeb and Pötscher, 2005; 2006; 2008).
Post-model-selection challenges follow that if acknowledged, often can be usefully addressed (Berk et al., 2013). We illustrate a form of post-model selection inference (PoSI) later using conformal prediction sets when we deliver provably valid uncertainty assessments applying nested conformal prediction sets to data restructured within a crossover design. Consistent with recent developments in statistics, design-based uncertainty estimates often accomplish what model-based uncertainty assessments cannot (Breiman, 2001b; Rubin, 2008; Rinaldo et al., 2019). We try to capitalize on design-based approaches in the analyses to follow.
3 AIRS Remotely Sensed Data
We provide next a relatively deep dive into the production of remotely sensed data from the Aqua spacecraft. Perhaps the primary message is that complicated hardware and software are required on which researchers must rely. There are at this point no “do-overs.” Moreover, the “Level 3” data fundamental to our analyses has the smallest spatial scale at the requisite geographical locations for remotely sensed data readily available in the public domain. Table 1 shows the variable downloaded for the analyses to follow.
The Aqua spacecraft responsible for the AIRS data was launched into polar orbit at an altitude of 705 km on May 4, 2002. Acquiring and processing AIRS data began in September of that year. Aqua crosses the equator during the ascending, or the northward daytime part of its orbit, at 1:30 pm local time. The descending, or southward, nighttime, part of its orbit crosses the equator at 1:30 am local time. Each half-orbit takes roughly 45 minutes, so local time of data acquisition is always within 22.5 minutes of equator crossing. AIRS successively scans across its 1500 km field of view taking data in 90 circular or elliptical footprints. As the spacecraft advances, the sensor resets and obtains another scan line. 135 scans are completed in six minutes, and this 90 by 135 footprint spatial array constitutes a “granule” of data. AIRS collects 240 granules per day, 120 on the daytime, ascending portions of orbits and 120 on nighttime, descending portions.
At each of the 12,150 footprints in a granule, AIRS observes Earth and its atmosphere in 2378 infrared spectral channels. Roughly speaking, the channels sense the surface and different altitudes in the atmosphere. The instrument counts photons at the different wave numbers (inverse wavelengths). These counts are converted to brightness temperatures ranging from zero to about 340 degrees Kelvin. Such data are called Level 1 products (Aumann et al., 2020). Certain atmospheric characteristics are related to photon emission, and these characteristics can be retrieved by solving complex sets of equations. The computational implementation is called a retrieval algorithm (Susskind et al., 2020), which converts spectra of brightness temperatures into estimates of geophysical variables. For us, the potentially relevant variables include surface temperature, tropopause height, and vertical profiles of air temperatures and water vapor mass-mixing ratios at various altitudes starting from the near-surface up to 250 mb pressure. These data are known as Level 2 products (Thrastarson, 2021).
The mission also provides a globally gridded version, summary product that averages Level 2 data on a one-degree-by-one degree daily spatial grid. These are Level 3 data products (Tian et al., 2020) intended for users who do not require the higher spatial and temporal resolution of Level 2. Typically studies of global climate processes use Level 3 because it is easier to compare with the output of climate models. Many variables are provided in the AIRS Level 3 daily product. The ones of interest to us for this study are shown in Table 1. Their variable names are “official” in that those are the names that come with the downloaded data. We shorten some names and change some names to facilitate data processing and the construction of tables and figures. But the mapping from the names in Table 1 to names used in subsequent tables and figures should be clear.111 In Table 1, the phrase “Pressure levels of temperature and water vapor profiles” means that atmospheric pressure is proxy measure for altitude for both temperature and the water vapor profile. The former is in degrees Kelvin. The latter is a lot like relative humidity at different altitudes but is the proportion of the total atmospheric profile that is water vapor at a given altitude. The total atmospheric profile is the proportional representation of different constituents of the atmosphere at a given altitude such as oxygen, carbon dioxide, nitrogen, and water vapor.
Name Dimensions Description
Latitude $360\times 180$ Latitude at the center if the grid cell
Longitude $360\times 180$ Longitude at the center of the grid cell
LandSeaMask $360\times 180$ 1 = land, 0 = sea
Topography $360\times 180$ Topography of the Earth in meters above the geoid
StdPressureLev $24\times 1$ Pressure levels of temperature and water vapor profiles
(we use only 1 – 12). The array order is from the surface
upward.
SurfAirTemp $360\times 180$ Temperature of the atmosphere at the Earth’s surface in
degrees Kelvin
TropHeight $360\times 180$ Height of the tropopause in meters above sea level
Temperature $360\times 180\times 24$ Temperature of the atmosphere in degrees Kelvin (we
use only bottom 12)
H2O MMR $360\times 180\times 12$ Water vapor mass-mixing ratio in grams/kilogram
Table 1: AIRS Level 3 variables used in this study, along with their dimensions and brief descriptions.
4 Research Design
The research design is influenced substantially by the rare events problem. We capitalize on crossover, quasi-experimental principles. The study units serve as their own controls. Forecasting skill likely is the beneficiary. They also provide a foundation for overcoming statistical problems from rare events.
To begin, any research design in service of forecasts of extreme heat waves must address the question “compared to what?” The obvious answer is compared to “non-heat waves.” But the details can be subtle. Summer months in many parts of the world can manifest several very warm days in a row that may be unusual, and when juxtaposed to recent history, may seem extreme.
We begin with the 2021 Pacific Northwest extreme heat wave affecting parts of the U.S. and Canada. During the last four days of June, the Pacific Northwest experienced what many observers called an extreme heat wave. Although there are debates over what exactly constitutes an extreme heat wave, this event seems to qualify from almost any informed perspective. With so many extant definitions (Perkins, 2015, Barriopedro et al., 2023), and nothing close to a consensus on best practices, we will provisionally operationalize an extreme heat wave as a single discrete high temperature event over several days under a slowly moving “heat dome” (Li et al., 2024). Other formulations certainly are possible, and perhaps even preferable, depending on the policy and/or research questions asked.222 The many possible definitions of a heat wave leads to some clumsy wording. The term “true” heat wave implies broad definitional agreement. The same applies to “actual” heat wave. The use of “extreme” heat wave can be quite vague unless “extreme” is defined. Even the term “heat wave” also can mean that we are using a widely accepted definition. We use all four at various times depending on context with the understanding that the definition just provided is what we intend to convey.
Figure 1: U.S. Pacific Northwest study region shown by the large red rectangle. Smaller rectangles are grid cells.
Media coverage of the June heat wave was by and large consistent with data from the Community Earth System Model (https:// www.cesm.ucar.edu/), which is not employed in the current analysis. All relevant June days were confirmed with the PNW AIRS data used in the analyses to follow. The same sources were used to identify the absence of extreme heat waves in manner described shortly.333 The spatial boundaries of an extreme heat wave can be difficult to determine. One complication is a heat dome responsible can cover many square kilometers and move slowly over time. Different fractions of potentially affected grid cells, nominally in the same locale, will fall under the dome from day to day. In the case of the PNW June 2021 heat wave, the four-day definition seems to be a reasonable approximation.
Figure 1 shows the location of our study area: a $12^{\circ}\times 12^{\circ}$ region over the northwest U.S. and southeastern Canada. For this analysis, we use daily data from these 144 grid cells, each grid cell indexed by location and day. The widely varying topography apparent from Figure 1 can be associated with very different microclimates that must be taken into account as we proceed. Clearly, surface temperatures can be quite dissimilar spatially whether or not there is an extreme heat wave. We will later consider similar data from the greater Phoenix area.
Figure 2: The extreme heat wave is compared separately to Faux Heat Waves 1 and 2. The top blue line shows the research design for the extreme heat wave. The middle blue line shows the research design Faux Heat Wave 1. The bottom blue line shows the research design for Faux Heat Wave 2. Each design has a pre-test, post-test, and predictors. The extreme heat wave is for June, 2021. Faux Heat Wave 1 is shifted back 4 Days, and Faux Heat Wave 2 is shifted ahead 30 days. The pre-tests and post-tests are each computed separately over grid cells as the surface temperature mean of 4 consecutive days.
To what should the PNW 2021 heat wave been compared? Our research design for the PNW AIRS data is shown in Figure 2: the top blue line for the extreme heat wave, the middle blue line for comparison “Faux Heat Wave 1,” and the bottom blue line for comparison “Faux Heat Wave 2.” The study units are the same 144 grid cells for the extreme heat wave and both faux heat waves. Much of the data for a grid cell can differ depending on the day on which remote sense data were collected. (See Table 1.)
Starting with the heat wave at the top representing the month of June, 2021, the post-test is the mean surface temperature for each grid cell over the 4 last days of the month. The pre-test is the mean surface temperature for each grid cell over the 4 sequential days immediately preceding the first post-test day. In both cases, means are used to help remove random variation in noisy measures of daily surface temperatures. Having a pre-test as well as a post-test confers benefits discussed in the next section. All of the predictor values are taken from 14 days before the beginning of the true heat wave. The same 144 grid cells constitute the post-test, pre-test, and lagged predictors.
Faux Heat Wave 1, shown as the middle blue line, has the same design configuration as the heat wave, but is shifted back in time 4 days before the extreme heat began. Its post-test does not include the extreme heat wave, which provides one answer to the question “compared to what?” Faux Heat Wave 2 at the bottom of Figure 2 also has the same design configuration as the extreme heat wave, but is shifted forward in time one month to July, 2021 such that the extreme June heat wave also is not included in the post-test. It provides as second answer to the question “compared to what?” For both faux heat waves, the same 144 grid cells constitute the post-test, pre-test, and lagged predictors.
Faux Heat Wave 1 has the advantage of a post-test falling immediately before the extreme heat wave post-test so that potential weather confounders might be quite similar. The risk is that some of the extreme heat wave characteristics may be in part a spillover from the immediately preceding several days; a larger temporal separation might be desirable. Therefore, Faux Heat Wave 2 provides a temporal separation of a month, but risks that summers in July may on the average differ from summers in June in ways confounded with surface temperatures.444 We also considered computing the pre-test using the grid cell means for the full set of June days before the heat wave. However, this approach builds in temporal warming trends in surface temperatures gradually unfolding during the June.
5 Surface Temperature Changes from Pre-Test to Post-Test
We transition to empirical analyses by separately examining the PNW heat wave summer temperatures and routine PNW summer temperatures. The response variable is gain scores of means grid cell by grid cell. For each grid cell, the gain score is the post-test mean temperature minus the pre-test mean temperature. It is just a usual difference in means that is computed separately for the actual heat wave and for each faux heat wave. Despite the name, gain scores can be positive or negatives and are sometimes called “change scores;” a gain score is nothing more than the change from pre-test to post-test. For the analyses immediately to follow, attention centers on the change for each grid cell in the mean surface temperature between the 4 day pre-test and the mean surface temperature during the 4 day post-test.555 Gain scores do not have to be a difference in means. They are more commonly a difference in scores, one for a pre-test and one for a post-test. For example, using a single temperature value the day before a heat wave and another single temperature value the day that heat wave begins would be an example. The length of their temporal separation would depend on the application.
Some grid cells are routinely warmer than others because of fix features such as urban heat islands. Gain scores control for fixed pre-test temperature differences between grid cells that might otherwise might be folded into post-test temperatures. Moreover, by removing earlier surface temperature variation that might be properly characterized as a nuisance in the study of heat waves, more accurate forecasts can result. These are a well-known properties of gain scores (Maris, 1998, Maxwell et al., 2018)
5.1 Gain Score Distributions for the Actual Heat Wave and Faux Heat Wave 1
Figure 3 shows the grid cell, surface temperature gain scores of means histogram for the extreme heat wave, for the Faux Heat Wave 1 and for their overlap. Variation in both sets of gain scores over grid cells is to be expected. The gain scores condition directly on pre-test surface temperatures and indirectly on fixed temperature-related features of grid cells such that dynamic forcings influencing surface temperatures can be better revealed.666 Statistical interactions between dynamic processes and fixed features of a grid cell certainly can occur and are not controlled by the pre-test conditioning. Estimating such relationships can be very challenging and is beyond the goals of this analysis, necessarily constrained by what the AIRS data measures.
Figure 3: fig: PNW Histograms of the gain scores of means for the extreme heat wave in red, faux heat wave 1 ind blue, and their overlap in purple.
Neither the actual heat wave histogram nor faux heat wave histogram has much resemblance to the canonical normal distribution. There is far too much mass toward the centers of both histograms. Nor is either distribution strongly skewed to the left or the right. There is nothing to suggest an extreme value distribution. The attempt to isolate the role of dynamic temperature process does not lead to popular “off-the shelf” distributions in gain scores.777 These are “eyeball” assessments that require some care because the aspect ratio of all three histograms produces rather flat-looking frequency profiles. When the histograms have an aspect ratio of 1 to 1 in height to width, the eyeball impression is close to a rectangle that is taller than it is wide. The 1 to 1 aspect ratio led to distorted displays when collected on a single page.
One might conclude that the two histograms are much alike. However, the extreme heat wave histogram has 3 of 120 grid cells with negative values and a gain score mean of 3.8. The faux heat wave distribution has 44 of 128 grid cells with negative values and a gain score means of 1.2.888 There are less than 144 grid cells in either histogram because when a dataset for each was built with all of the predictors, there are 20 or more observations with missing data. One can explore further how the distributions compare by plotting the two histograms in the same figure. The bottom display shows the distributional offset in more detail. The red bars are taken from the histogram for the extreme heat wave gain scores. The blue bars are taken from the histogram for the faux heat wave 1 gain scores. The purple bars show the overlap. Clearly, the two distributions are quite different.
More is going on than a simple distributional shift to the right. The distributional shape changes as well. The extreme heat wave histogram has more of its mass above its mean flattened and extended toward higher gain score values. Nevertheless, there is a substantial overlap between the two distributions. Although extreme heat waves can be very consequential, a substantial fraction of the gain scores of means are much the same for both histograms.
In the interest of space, Faux Heat Wave 2 is not yet considered. But Its histogram looks a lot like the histogram for Faux Heat Wave 1. We make more use of Faux Heat Wave 2 later to replicate some important findings for Faux Heat Wave 1. The key point now is that when examining histograms, the role of a 4 day lag versus a 30 day lead does not seem to make much empirical difference in the gain score histograms.
6 Fitting The Gain Scores of Means With AIRS Predictors
Differences between the two sets surface temperature gain scores have been described. As a step toward forecasting heat waves, one might wonder if the actual heat wave and Faux Heat Wave 1 have precursors that differ as well. Those precursors were shown in Table reftab: variables. To proceed, however, we need first to consider a bit further how the rare events problem might be mitigated.
6.1 Mitigating The Rare Events Problem
We have emphasized that extreme heat waves are statistically rare. But how rare depends on the population of which they are a part. For example, if one considers all days from 2017 through 2021 in the Pacific Northwest as a population of interest, there would be one extreme heat wave. Labeling all four of the last days in June of 2021 as heat wave days, there would be 4 extreme heat wave days for more than 1800 days overall (i.e., $365\times 5$). If one never predicted an extreme heat wave over all of those days, that prediction would be correct about 99.8% of the time without using any predictors at all. It would be, therefore, almost impossible to uncover the roles of predictors because fitting is already close to perfect. As noted earlier, one is stuck with very accurate and very trivial forecasts. Even if the data were limited to all of the days in 2021, a forecast of no heat wave would be correct 98.9% of the time. The fitting problems are nearly as catastrophic.
There is more promise using data only for the 30 days in June of 2021. The trivial forecast then will be right “only” 86.7% of the time. We can make further progress by using grid cells as the study units for the post-test, pre-test, and on the day two weeks earlier when the predictors are measured. Figure 2 shows the result for the true heat wave, Faux Heat Wave 1, and Faux Heat Wave 2.
The point of Figure 2 is solve the rare event problem across the three settings. If the actual heat wave scenario is compared to either of the two faux heat waves, there are in principle 144 true heat wave observations and 144 faux heat wave observations. If the extreme heat wave setting is compared to a pooled version of the two faux heat waves, there remain theoretically 144 observations in the heat wave scenario and then, 288 observations in the combined faux heat wave scenarios. In either case, the PNW extreme heat wave is not rare.
We are capitalizing on a variant of crossover, quasi-experimental designs (Lui, 2016; Maxwell et al., 2017). Crossover designs are common, especially in biomedical research and academic psychology. From Figure 2, all grid cells are exposed to extreme heat wave precursors at some specified time and the same grid cells are exposed to routine summer conditions at some other specified time. Because the grid cells serve as their own controls, subsequent forecasting accuracy is likely to improve.
However, there is a price to pay for these design adjustments to the rare events problem. Perhaps most important, we use a form endogenous sampling when the three design configurations in Figure 2 are imposed to all grid cells. It is unlikely that any study of rare extreme heat waves could be undertaken in a naturally-occurring population such that one-third of the data were found in each of three different study scenarios. Forcing the data to be balanced in this manner can produce misleading statistical results (Manski and Lerman, 1977; Manski and McFadden, 1981). We will return to this problem later to consider several possible solutions. For now we continue to descriptively unpack how extreme heat wave surface temperatures can differ from faux heat wave surface temperatures.
6.2 An Application of Supervised Learning
We deploy a supervised learning algorithm that can have good forecasting skill. Just as shown in Figure 2, the response variable is grid cell gain scores of means for the true PNW heat wave and Faux Heat Wave 1. (Faux Heat Wave 2 will be used later.) The predictor values are taken from two weeks earlier for all variables shown in Table 1 except ”SurfAirTemp”, which is used to construct the pre-test and post-test. There are 286 observations, 144 grid cells each from the extreme and faux heat wave less two grid cells having missing data.999 Two grid cells were deleted because of missing data. We considered two forms of widely used and successful supervised learning: random forests (Breiman, 2001a) and stochastic gradient boosting (Friedman, 2001). Both can be seen as variants of nonparametric regression, are very easy to train, and often the default tuning parameter settings work well. A complication is that there is strong dependence between many of the AIRS predictors. For example, the measured temperatures at different altitudes are often correlated at values above .95.
Random forests handles strongly correlated predictors better than stochastic gradient boosting because it is comprised of a large ensemble of regression or classification trees with the partitioning for each tree undertaken using a small random subsets of predictors. Sampling predictors (without replacement) for each split can dramatically reduce the impact of high multicollinearity. We settled on the implementation of random forests in R, ported by Andy Law and Matthew Wiener (Liaw and Weiner, 2002) from the original Fortran by Breiman and Cutler (Breiman, 2001a), which includes several important auxiliary algorithms that can dramatically increase algorithmic transparency.101010 We briefly considered using deep neural networks. But the vastly different units for some of the predictors caused problems for the back propagation. Dependencies between the predictors also caused computational difficulties. Once both problems were addressed, and tuning was able to proceed, the fitting performance was a bit worse than for random forests and gradient boosting. Also many of the auxiliary procedures that shine some light into the algorithm black box (Molnar, 2022) were no longer readily available.
In this highly exploratory phase of the work, we were motivated by possible differences between how the same predictors performed depending on whether the data came from the true extreme heat wave or a faux heat wave. Work on quasiresonant amplification is a provocative illustration (Petoukov et al., 2012; Mann et al., 2018; Li et al., 2024). In response, we initially applied random forests separately to the extreme heat wave dataset and to the faux heat wave dataset structured as in faux heat wave 1. We computed 500 regression trees for each. About 75% of the gain score variance was accounted for using either the true heat wave data or the faux heat wave data. By the MSE criterion, fitting performance was promising and much the same for both.
But all else was quite different. For example, the fitted values for the heat wave histogram were farther to the right than the fitted values histogram for the faux heat wave, which follows in part from the different gain score distributions discussed earlier. Perhaps most telling, different predictors dominated the two different fits depending on whether the response variable was the gain scores for the extreme heat wave or the gain scores for the faux heat wave.
Figure 4: Variable importance plot for the random forest fit of the extreme heat wave (top) and Faux Heat Wave 1 (bottom). For both, the five more important proctors are shown. The variable names were shortened for ease of writing code but are more fully explained in the text.
Figure 4 is the variable importance plot for the extreme heat wave on the top and Faux Heat Wave 1 on the bottom. The five most important predictors for each are shown. The 14 day lagged predictors are listed in the left margin. Predictor importance is greater for predictors higher in listing. The metric at the bottom is the percentage increase in the gain score mean squared error (MSE) when each variable by itself is randomly shuffled such that associations with the gain scores and with the other predictors are driven toward zero. The increase in MSE measures the decrement in fitting performance.
Starting with the top figure at the top of variable list, when a grid cell’s topography (meters above sea level) is shuffled, the MSE increases by about 3%. If a grid cell’s land/sea indicator is shuffled, the MSE increases by about 1.7%. If two temperatures at relatively high altitudes are shuffled, MSE increases about 1.5% (i.e., 1% + .5%). If water vapor mass-mixing ratios in a middle altitude is shuffled, the MSE increases by about .5%. Although these are for the five largest increases in fitting error, in numerical terms they are quite small relative to the overall variance accounted for by the random forest. A significant reason is that the predictor contributions are shared with other predictors due to interaction effects that are not included in these MSE increases. This explains why the fixed grid cell features of elevation above sea level and whether grid cell is over land or over water appear in the figure.111111 It is possible to show these effects (Molnar, 2022), but the number of predictor pairs, triplets, and more that would need to be examined make this strategy computationally impractical. Interactions between predictors are included in the overall fit even though we cannot separately identify and characterize them.
The bottom display in Figure 4 is the variable importance plot for the faux heat wave. The two fixed predictors not shown because they have very small contributions to the fit. What matters most are several measures of temperature at different altitudes 14 days earlier. Grid cells that have larger gain scores during the extreme heat wave are warmer 14 days earlier as well. There is some form of temperature temporal dependence. But the point is that despite comparable fitting performance, the five predictors that uniquely matter most for the fit appear to differ depending on whether random forests is applied to the true heat wave data or the faux heat wave data.
In summary, both variable importance plots show which predictors are driving most the random forests’ fitted values. The predominant predictors differ depending on whether the gains scores belong to the actual heat wave or to the faux heat wave. This also is consistent with the claim that the extreme summer heat is not just an unusually harsh version of ordinary summer temperatures.
However, there is substantial dependence among the AIRS predictors. We repeated the random forest fitting several time for both the extreme heat wave and faux heat wave. The modest amount of randomness built into random forests sometimes led to disconcerting changes in the most important predictors identified. This instability motivated a look at faux heat wave 2. One can apply the exact same design used earlier but replace faux heat wave 1 with faux heat wave 2 and apply random forest as before. The crossover design now separates the actual and faux heat wave by 30 days instead of 4 days. Consequently the risks from spillover contamination are dramatically reduced. But the large gap in time can allow for differences in the grid cell data making comparisons between the predictors of the actual and faux heat waves more complicated. For example, on the average, weather conditions in June and July related to surface temperatures in may differ. There can be less rain and lower soil moisture in July, for instance. According to NOAA National Centers of Environmental Information, 1.9 inches of rain fell in Seattle in June of 2021, and 0.0 inches of rain fell in Seattle in July of 2021 (https://www.ngdc.noaa.gov/).121212 In the winter months, the rainfall in Seattle is much more substantial.
Despite such possibilities, the random forest fit to the faux heat wave gain scores of means in July is again quite good. About 70% of the variance in means-based gain scores was accounted for. Variable importance calculations determined the top five predictors as temperatures at middle to high altitudes, just as for the earlier faux wave. No water vapor mixing ratio made the cut. These results are similar to those for faux heat wave 1, despite some differences in the research design. Perhaps neither the difference in pre-test to post-test separation time nor the differences in weather patterns between June and July matter much. And there is some additional evidence that extreme summer heat waves may differ structurally for ordinary summer heat. Yet, the substantial dependence between predictors remains worrisome. There is ample reason to consider further the predictors that appear to matter most, which also tend to differentiate the actual June heat wave from the faux heat waves in June and July. Another data analysis approach may be needed.
7 Applying Genetic Algorithms To The Fitted AIRS Data
We have shown that the distribution of gain scores from the actual extreme heat wave and from Faux Heat Wave 1 differ in important ways. There also is some evidence that for the actual and faux heat waves, precursor variables 14 days earlier can differ as well. But dependence between the 14 day lagged predictors makes claims about such differences somewhat tenuous.
The statistics literature has long suggested ways, inconclusively, to think about and analyze variables that are highly correlated (Alin, 2010). Some recent work seems especially relevant for our setting: variable selection combined machine learning (Chan et al., 2022). But all approaches we have seen to date accept the data on hand as the venue for progress. The results are a long way from closure.
If there were such a thing as an ideal extreme heat wave and such a thing as an ideal set of warm summer days, perhaps the empirical messiness of heat wave drivers could be reduced substantially. Recall that the June PNW faux heat wave had many gain scores similar to those of the June actual heat wave. This is not surprising because surface temperatures are produced by normal weather and topographical forcings likely combined with dynamic forcings when a heat dome arrives. Ubiquitous factors such as earlier surface temperatures are still in play. Perhaps for similar reasons, many of the predictors available in the AIRS data must be highly correlated making it extremely difficult to isolate their separate effects. Might abstractions from extreme heat waves and usual summer days help? Under the right circumstances perhaps an ideal extreme heat wave might have only a few dominant predictors and an ideal garden-variety series of very warm days also might have only a few, but different, dominant predictors.
We can borrow from the social sciences the concept of an “idea type” or a “pure type” (Garth and Mills, 1958) that abstracts the essential features of some phenomenon much like many key concepts in climate science. Then, perhaps a genetic algorithm can be employed to construct two ideal types: a synthetic population of extreme heat wave grid cells and a synthetic grid cell population of normal sequences of very warm summer days. Comparisons between the two might allow for a more definitive reading of their similarities and differences.
A random forest was trained as before using 500 regression trees for the June extreme heat wave. The gain score data used for training were the same as the gain score data employed in the earlier analyses. A trained and averaged ensemble of regression trees can serve as a survival function for a genetic algorithm’s construction of an ideal, synthetic population of grid cells experiencing a true heat wave.
In the same manner, a random forest was trained as before using 500 regression trees on June actual heat wave 1. The gain score data were the same as used in the earlier analyses. A trained and averaged ensemble of regression trees can serve as a different survival function for a genetic algorithm’s construction of an ideal, synthetic population of grid cells experiencing a faux heat wave.
The two survival functions were used as input to the procedure GA in R written by Luca Scrucca (2013). The standard defaults for numeric predictors worked well. For both survival functions, 5000 iterations were specified leading to 5000 synthetic populations with 100 grid cells each having increasing fitness on the average. As a first approximation, one can think of a genetic algorithm as jittering predictor values, computing new fitted values from the survival function using the jittered predictors, saving the subset of the grid cells with the larger gain scores, and repeating this process many times until there is a population that includes only very fit grid cells (i.e., having large gain scores).131313 The GA software in R has several tuning parameters with good default values.(1) For each synthetic population of grid cells, the predictors from the five grid cells with the highest computed surface temperatures were not altered. (2) The crossover probability for each possible pair of predictors was set at 0.8. A crossover refers to a pair of “parent” grid cells combining their predictor values to form an “offspring” grid cell. (3) Because all of the predictors are treated as numeric in this application, the predictor values for an offspring grid cell are computed as the average of the corresponding predictors from a random pair of parent grid cells. (4) The mutation probability for each grid cell’s predictor was set at 0.1. (5) The mutation method used replaced the value of the existing numeric variable with a random value within a specified range drawn at random from a uniform distribution. The size of a synthetic population is not a tuning parameter but 100 grid cells seemed a good compromise between an ability to examine the output visually in detail and having a sufficient number of observations to arrive at sufficiently stable parameter estimates.
The grid cell mean gain score for the final synthetic population produced by the actual heat wave survival function was 7.26 degrees K. The grid cell mean gain score for the final synthetic population produced by the faux heat wave 1 survival function was 5.76 degrees K. Because of differences in the two survival functions, the genetic algorithm could not equate the two. One possible inference, already mentioned, is that the physics for a true heat wave may differ somewhat from the physics of a faux heat wave.141414 Predictor dependence was reduced substantially because of the predictor randomness the genetic algorithm introduced.
Table 2: Genetic algorithm predictor “solution” values for true heat wave and the faux heat wave 1.
Predictor True Heat Wave Faux Heat Wave
metersabove 623.02 395.78
land 0.73 0.67
temp4 274.48 228.16
temp5 268.84 232.32
temp6 259.13 224.15
temp7 246.98 225.02
temp8 233.93 211.24
temp9 215.74 245.85
temp10 217.13 260.26
temp11 218.14 227.19
temp12 220.84 265.94
mix4 5.33 6.97
mix5 3.44 1.47
mix6 1.67 0.76
mix7 0.61 0.16
mix8 2.60 6.86
mix9 1.31 5.84
mix10 4.47 3.90
mix11 7.68 5.73
mix12 2.84 5.20
tropheight 11766.76 6728.78
Table 2 provides a bit more evidence about the ways in which an ideal true heat wave can differ from an ideal faux heat wave. Shown are the predictor values the algorithm fashioned in the final iteration that produced grid cells with superb fitness in the metric of surface temperature gain scores. The predictor values for the extreme heat wave are in the left column. The predictor values for faux heat wave 1 are in the right column. Comparisons between the two columns can be used for variable selection in a simple application of feature engineering.151515 This is an example of cherry picking that can invalidate conventional statistical inference. However, the analyses to follow are able sidestep this problem.
Compared to the faux ideal heat wave, the true ideal heat wave is characterized by substantially higher grid cell, gain score elevation (i.e. 623 meters versus 396 meters) and higher grid cell gain score atmospheric pressure at the tropopause (i.e. 11,767 hPa versus 6,729 hPa). The ideal faux heat wave is characterized by higher grid cell, gain score atmospheric temperatures closer to the ground and lower grid cell, gain score atmospheric temperatures at high altitudes. The role of the atmospheric water vapor mass mixing ratios is inconsistent. There is now perhaps a bit better sense of which predictors are related most strongly to a extreme heat waves compared to a faux heat waves.
8 Forecasting Extreme Waves Using a Binary Outcome
We have arrived at a pivotal research destination. If the forecasting goal is help inform policy options two weeks (or more) before an extreme heat wave arrives, one must anticipate how forecasting could undertaken in practice. To be useful, the predictors must be measured and readily available well in advance of a possible heat wave. This implies ongoing data collection on a regular basis.161616 There are also second order requirements. For example, the data collection requires continuity in naming conventions, which factors measured, the temporal and spatial scales, the measurement technology used, and data file formats. If any of these change, forecasting skill can be badly damaged. At the very least, abrupt variation in the data over time will likely be introduced. In addition, promising predictors should not be constant over time. For each grid cell, altitude above sea level, for instance, is a constant that will not change whether or not an extreme heat wave subsequently occurs. It has no forecasting capacity.
Consider again briefly the operationalization an extreme heat wave discussed earlier. An extreme heat wave is a discrete event; it is either present or not present. The task becomes forecasting a categorical binary outcome. Moreover, when using earlier temperature gain scores, we were fitting a consequence extreme heat wave precursors and perhaps the impact of a heat dome. We were not explicitly forecasting a discrete, extreme heat wave.
Returning to the design in Figure 2 for the June true extreme heat wave, we drop the pre-test and recode the post-test to equal “1.” For purposes of comparison, we use July, 2021 as the setting for a faux heat wave (i.e, faux heat wave 2). The pre-test is dropped and the post-test is recoded to equal “0.” The same 144 grid cells are used for a new extreme heat wave design and for the new faux heat wave design; the same grid cells are exposed to one climate regime in June and then another climate regime in July. The June data are then “stacked” above (or below) the July data to form a single dataset. We have an explicit crossover design including both an extreme heat wave and a faux heat wave. The classification format is balanced, and the rare event problem seems resolved.
But the rationale goes a bit deeper. When the data are stacked, each study unit once again serves as its own control. There is no longer a need to compute a pre-test score. We use the last 4 days of July for the faux heat wave because the time offset between the actual and fax heat wave is now 30 days. The chance of spillover effects becomes very small. This helps justify the absence of a pre-test. Because the earlier results for Faux Heat Wave 1 and Faux Heat Wave 2 were so similar, the different seasonal conditions between June and July probably do not matter much for our classification analysis. Finally, by replacing a regression random forest with a classification random forest, we set the stage for latter of estimating forecasting uncertainty with conformal prediction sets (i.e., not regions).
Random forests was retrained as a classifier using the new data structure. All of the predictors used previously were candidates for the classification exercise. As a start, however, we selected three predictors from Table 2 that seemed most promising: (1) atmospheric pressure at the tropopause, (2) atmospheric water vapor mass mixing ratio at level 8, and (3) the temperature at level 8. For each, we considered the gaps shown in Table 2 between values in the actual heat wave column and values in the faux heat wave column. For the mixture and temperature variables, other proximate altitudes could have been selected instead because proximate altitudes tend to be substantially correlated. Dependence between the selected three predictors by themselves was not a serious complication.
No causal claims are being made. Our earlier analyses are consistent with the expectation that relatively few predictors can suffice for promising classification accuracy. High classification accuracy can be a precursor to high forecasting accuracy.171717 Classification accuracy is not the same as forecasting accuracy. For classification, the outcomes of interest are known and included in the data. Using a statistical classifier, the goal is to correctly identify each known outcome. For forecasting, the outcomes are unknown. If they were known, there would be no need to forecast them. The goal is to correctly project what the outcomes will be. Although as a statistical matter, high classification accuracy can support high forecasting accuracy, forecasting introduces additional assumptions and uncertainties that must be properly addressed. But high classification accuracy does not necessarily imply credible causal inference.
Our random forests results are encouraging. The overall classification error rate is 5.0%. In this setting, a false positive can be defined as incorrectly classifying the faux heat wave as the true heat wave. The false positive rate is 5.7%. In this setting, a false negative can be defined as incorrectly classifying the true heat wave as the faux heat wave. The false negative rate is about 4.1%. In short, the random forest classifier almost always correctly identifies the true positives and the true negatives. This can enormously benefit conformal prediction inference.
Figure 5: Variable importance for accurate classification of heat waves. The variable names were shortened for ease of writing code but are more fully explained in the text.
Figure 6: Partial Dependence Plots for the PNW Forecasts.
Figure 5 is the variable importance plot from random forests. The metric at the bottom of the figure is the reduction in classification accuracy average over both outcome classes when each variable in turn is randomly shuffled with the other variables fixed at their empirical existing values. For example, classification accuracy declines by nearly 30 percentage points if temperature at altitude 8 (only) is shuffled. In short, Figure 5 shows all three predictors are important for classification accuracy.181818 Average classification accuracy is used because the false negative and false positive rates are very similar. There is at this point no need to focus on either outcome class by itself.
Figure 6 show the partial dependence plots for the three predictors. A partial dependence plot displays the conditional mean of the fitted values for different values of a given predictor, all other predictors held constant at their mean values in the data (Friedman, 2001; Greenwell, 2017). These values necessarily are unchanging as the selected predictor’s values change; they are fixed. In this way, the possible impact of correlated predictors is controlled, but not with conventional covariance adjustments used parametric classification (e.g., logistic regression).191919 Unlike covariance adjustments, no assumptions need be made about how the predictors are related to one another. The vertical axis in Figure 6 is in probability units with black dots the fitted values. The blue line overlaid is a lowess smoother (Cleveland and Devlin, 1988).
The partial dependence plot at the top is for the predictor atmospheric temperature at altitude 8. Judging by the smoothed fitted values, there is an S-Shaped relationship with the probability of a heat wave but the fitted values are actually flat until the temperature is approximately 232 K and after it reaches about 235 K. Sparse data at in the tails of the temperature distribution may be the explanation. Overall, the relationship is positive, and ranges from near 0.0 to near 1.0 with temperature increases two weeks earlier.
The partial dependence plot in the middle is for the atmospheric pressure at the tropopause. Judging by the smoothed fitted values, there is as atmospheric pressure rises nearly a straight line increase in the probability of a heat wave from about a 0.0 probability to almost a 1.0 probability. However, fitted values are flat until atmospheric pressure reaching 11500 hpa, perhaps an impact of sparse data for lower atmospheric pressures.
The bottom partial dependence plot shows the relationship between heat wave probabilities and the atmospheric water vapor mass mixing ratio at altitude 8. Consistent with its weak importance to the fit, the relationship shown by the smoother is seems haphazard. The fitted values appear disorderly. However, the disappointing performance does not seem to undermine the overall excellent classification accuracy
8.1 Forecasting Uncertainty and Conformal Prediction Sets
Given the data available and possible peculiarities of PNW grid cells, classification accuracy is very promising. A next task is to turn classification into forecasting coupled with valid estimate of uncertainty. We use nested conformal prediction sets. For an unlabeled grid cell, conformal prediction sets can produce an extreme heat wave and/or a no extreme heat wave forecast that is provably correct at a predetermined probability. Because the random forest classifier discussed earlier was so accurate, there is good reason to expect that such forecasts will be very accurate as well. To address uncertainty in practical manner, we proceed with nested conformal prediction sets under the popular “split sample” approach (Kuchibhotla and Berk, 2023).
The key assumption is that the dataset used is comprised of exchangeable observations. This means that the order in which the observations are realized does not matter for the joint probability distribution of the data. When forecasts are needed for new cases without labels, the new observations also are realized in an exchangeable manner from the same population as the data on hand.
Our crossover design buttresses the exchangeable assumption. It does not matter whether the data from June were realized in July and data from July were realized in June. Their joint probability distribution would be the same. Recall that because of the 30 day disparity between the June actual heat wave and July faux heat wave, spillover effects are highly unlikely, and there is no other apparent connection in the data between the late heat wave events in June and the late faux heat wave events in July.202020 The issues can be subtle. This means that if for the data on hand the designated months were swapped, the histogram for the full dataset would be unchanged. It does not mean that grid cells measured in different months are exposed to the same climate conditions.
As the split sample approach requires, the data on hand were randomly split into two disjoint samples, one used to train a classifier, and one used to construct conformal prediction sets. Some call the second split a “test” sample. Some call it “calibration” data. One advantage of this approach is that most any classifier can be used, even methods like stepwise logistic regression, which aggressively capitalize on data snooping that can invalidate convention statistical inference (Berk et al., 2013). Such problems and others are not transported from the training data to the calibration data because the calibration dataset only gets to work with the classifier output, not the processes by which the output was produced. The calibration data distribution provides a baseline used to evaluate hypothetical forecasted outcomes for unlabeled data for which a forecast is sought.212121 Technically, only the calibration data need to be exchangeable.
Nested conformal prediction sets (Gupta et al., 2022) was applied to the crossover data for June and July of 2021. A predetermined coverage probability of .75 was specified. When for a given, unlabeled, grid cell a claim is made that the true future outcome is included in the prediction set, it will be correct at least 75% of the time; the odds in favor of the truth of that grid cell’s forecast are at least 3 to 1.
For illustrative purposes, we used 10 grid cells from the training data as if their true outcome class was unknown. For each grid cell, there was a single member in the prediction set, which represents ideal precision. Based on the outcome classes in the calibration data, the forecasts for all 10 cases were correct. This is no surprise because the random forests classifier identified the outcome classes so well. In retrospect, a more demanding coverage probability than .75 could have used (e.g. .90) because the random forests classifier was so accurate.
9 A Replication Study in Phoenix, Arizona
There are at least three reasons to undertake a replication in a new setting. First, the forecasting results just reported capitalize on extensive data exploration that can lead to cherry picked results; post-model-selection inference can be concern regardless of our efforts made to respond to the problem. Second, our findings acquire far more credibility and usefulness if they apply to areas in addition to the Pacific Northwest. That credibility and usefulness can affect subject-matter considerations and pressing policy concerns. Finally, other geographical areas bring not just different climate scenarios, but different data and different data problems. Especially important is to determine if our results are robust to new data flaws and research design constraints.
To address these matters of generalizability, we obtained AIRS daily data for the Phoenix area for 63 grid cells from 2019 through 2023. Data were not available for 2024. The measured variables were the very same used in the PNW analysis. However, the topography in the greater Phoenix area is not just very different from the PNW, but far more homogeneous. One might wonder, for instance, if a quasiresonant explanation applies.
We began with a simple and ambitious aspiration. Using a crossover design much like before, training with random forests on the same variables used previously, and with similar algorithmic methods to that used for the PNW forecasts, could we reproduce the PNW results in Phoenix? An exact replication would be effectively impossible to produce, but could we get results properly to inform affirmative judgments about reproducibility?
There were immediate and obvious design issues. We had about half the number of grid cells for the Phoenix area than for the Pacific Northwest. There was also a substantial number of grid cells with missing data. Between the two, we anticipated a much smaller number of observations overall. One result could much weaker power and even more sparsity. We needed a way to increase the number of usable grid cells within in the intent of our earlier crossover design.
The Phoenix heat dome arrived at the beginning of July 2023, not at end of June, 2021, as it had in the Pacific Northwest. For an alternative crossover design, we proceeded the much as before, but the faux heat waves A and B for ordinary summer temperatures were for July of 2021 and July of 2022 respectively, before the actual Phoenix extreme heat wave. This means that the predictors lagged 14 days earlier would fall toward the middle of June for 2021, 2022 and 2023. For all three Phoenix years, we used a binary response variable coded “1” for the heat dome in July of 2023, and “0” for no heat dome in July of 2021 and July 2022.222222 We could have included a greater number of years back in time, but all would for faux July heat waves. The number of appearances made by each grid cell could be increased along with the overall number of observations. But we were concerned that the binary response variable would become badly unbalanced.
The coding is consistent with that done for the PNW forecasting analysis. All three time periods were then stacked to make single data set; each grid cell in principle was observed in three different years and in principle there would be 189 observations representing one extreme heat wave and two faux heat heat waves. The crossover design would again provide sufficient control for temporally stable features of grid cells.
Figure 7: The True Heat Wave is compared simultaneously both to Faux Heat Waves A and B. Each has a binary post-test, and predictors 14 days earlier. The True Heat Wave began the first day July, 2023. The two faux heat waves began on the first day of July 2021 and the first day of July 2022. Were we trying to be fully consistent with notation used in the research design for the PNW, the time indices start at t0 on 6/1/2021, and the time index values for the true heat wave and the 2022 faux heat wave would be much larger.
But the potential problem from missing data remained. When we removed all rows (i.e., grid cells) with any missing data, we lost 186 grid cells from the total 189. Clearly this was disastrous. Fortunately, when the included variables were limited to just those selected for the PNW forecasts, the number of grid cells fell to “only” 128. Because the data were lost through cloud cover and occasional downtime on the remote sensing satellite, one can argue that the data were missing in blocks conditionally at random, and exchangeability still holds. Nevertheless, the large amount of missing data raises concerns about the representativeness of the Phoenix data and about some instabilities that could be introduced. The design we used is shown in Figure 7.
Figure 8: Partial Dependence Plots for Phoenix Forecasts
We were pleasantly surprised. The error rate for Phoenix was a little under 4%, roughly comparable to the error rate for the PNW. As before, all three predictors made important contributions to the fit although their rank order differed. The atmospheric water vapor mass mixing ratio was now the top ranked. A different importance ordering was anticipated because the topography and climate in the Phoenix area is so very different from the topography and climate in the Pacific Northwest.
Despite great differences in climate and topography, Figure 6 for the PNW and Figure 8 for the greater Phoenix area have their smoothers showing much the same S-Shaped relationship for atmospheric temperature at an altitude of 8. Both relationships begin flat for the lowest values before the slope turns positive, and end flat at the highest values. The Phoenix relationship is shifted a bit to the right in response to its higher temperatures. It also somewhat less smooth because of data that are more sparse.
Comparing the PNW and Phoenix smooths for tropopause height, both start out flat before turning positive. For the PNW, the overall pattern is roughly linear and positive thereafter. For the greater phoenix area, the relationship turns negative at around 15,500 meters.
For the water vapor mixing ratio, neither the smoother for the PNW nor for the greater Phoenix area show easily interpreted relationships. This helps explain why in both geographical settings, the water vapor mix contributes little to the fit. The good news is that it seems not to be needed. In both settings the fit is excellent, and the two variables largely responsible have somewhat similar relationships with the presence of an extreme heat wave. Finally, we again employed nested conformal prediction sets. As before, forecasts were very precise (i.e., one outcome class chosen) even with coverage probabilities of .90.
It is important to emphasize that while the PNW analysis was steeped in extensive data exploration, the Phoenix analysis had none. There was no cherry picking of results whatsoever. Yet classification accuracy was the nearly the same for both and two of the three partial dependence plots had much the same look and feel despite some differences in detail.
10 Crossover Design Balance and Problems from Endogenous Sampling
Recall that the crossover designs for the PNW and for the greater Phoenix area began by constructing the data for a true and a faux heat waves with the same grid cells in each; each had the same number of observations. This led to excellent balance that remedied the rare events problem. However, it is challenging to think of a naturally occurring population of grid cells in which such balance could have occurred. For example, if the population were comprised of all grid cells in the pacific northwest for each day during June of 2021, for about 13% of the days, a grid cell would be subject to an extreme heat wave. Yet, in our designs, for approximately 33% or 50% of the days, an actual heat wave would have produced extreme surface temperatures. The marginal distribution of the curated 0/1 response variable was no longer distributed like the 0/1 response variable for the month of June. Our curation process oversampled extreme heat wave days, which is a form of endogenous sampling (because the response variable is treated as endogenous).
As noted earlier, there is a literature on problems that can result when data collection alters the prior distribution of the response variable compared to its distribution in the population of interest (Manski and Lerman, 1977; Manski and McFadden, 1981). For example, one might over-sample cancer cases that would otherwise be too small a fraction of the data for sufficient pattern recognition accuracy (Wang et al., 2021). But the price is empirical results that can lack many desirable statistical properties. There can be, for instance, badly distorted estimates of fitted probabilities.
For discrete response variables, several remedies have been proposed (Cosslett, 1993; Waldman, 2000). An entire dataset can be weighted in one of several ways so that the endogenous (response) variable once again can provide a consistent estimate of the response variable’s distribution in the target population. Sometimes the weighting is readily apparent. Sometimes it is buried in the statistical machinery along with performance enhancements.
Appropriate weighting is easily undertaken.232323 We take a transparent approach because we have seen no formal proofs of the more elaborate proposed remedies for random forests, gradients boosting, or nonparametric classifiers more generally, when used for forecasting. For example, whether or not there is an extreme heat wave is the endogenous (response) variable in Figure 7. Suppose $P(heat)$ is the probability of an extreme heat wave in the target population. $P^{\ast}(heat)$ is the probability of an extreme heat wave in the curated data on hand. The two probabilities are different because of the manner in which the data collected and organized. There is a pair of remedial weights (Manski and Lerman, 1977; Manski and McFadden, 1981). For each grid cell $i$ that experienced an extreme heat wave, $weight_{heat,i}=\frac{P(heat_{i}=1)}{P^{\ast}(heat_{i}=1)}$. For each grid cell $i$ that did not experience an extreme heat wave, $weight_{noheat,i}=\frac{P(heat_{i}=0)}{P^{\ast}(heat_{i}=0)}$. These weights can be incorporated easily within a weighting option commonly available in machine learning software.242424 As before, a heat wave is coded equal to “1.” The absence of a heat wave is coded equal to “0.”
There are several difficulties with such solutions. First, in many applications it is not apparent what the proper population is or even should be. An earlier example was whether the population for studying a rare extreme heat wave should be daily ARIS data for the most recent 5 years, daily AIRS data for the most recent year, or daily AIRS data for the most recent month. The choice can depend heavily how heat wave forecasting will be undertaken in practice. If essential responses to extreme heat waves are determined month to month, monthly data might be preferred. But then, a trained forecasting algorithm for each month would be needed, perhaps updated year to year. The work load could be reduced if forecasters were prepared to bet that rare extreme head waves only occur in the summer months. The forecasters would then be backing into the kinds of designs we used in the PNW and greater Phoenix area.
Second, for rare discrete events, the proper weighting proposals often recommended resurrect the very problem that the oversampling of rare events tried to remedy; it is not clear that the proposed remedies anticipated forecasting very rare events. The oversampled rare events are made empirically rare once again and preferred machine learning software may not run at all, or may run but give consequential warming messages. For example, the fitted probabilities may fall outside of the 0 to 1 range.
Third, Therneau and Atkinson (2023) have made clear that for tree-based classifiers such as classification trees, the prior distribution of the response is related to the ratio of the number of false negatives to the number of false positives (or the reciprocal). That ratio is often used to represent the relative costs of false negatives to false positive (i.e., all costs, not just monetized costs). For example, if for a binary response variable, a classifier produces 40 false positives and 20 false negatives, there are 2 false positives for every false negative. This can be interpreted to mean that false negatives are twice as costly as false positives, which may help explain why the classifier fits the data so that there are fewer of them; one false negative is “worth” two false positives.
More fundamentally, because the false negatives and false positives can have different costs, the algorithm is minimizing risk, not loss, a formulation convincingly introduced in the classic book by Luce and Raiffa (1957). It is no longer clear how endogenous sampling problems apply because the prior distribution of the response variable and differential losses for false negatives and false positives are related, and together they affect the risk. The underlying statistical machinery might be more appropriately built on asymmetric risk functions applicable to algorithmic forecasting (Ridnick et al., 2021). In fact, the entire discussion of unbalanced classification problems might be usefully reframed to make the consequences of false negatives and false positives a matter for asymmetric loss functions to address.252525 The ratio of false negatives to false positives is something about which stakeholders might care deeply. How much more (or less) costly a failure to correctly forecast an extreme heat wave (i.e., a false negative) may be compared to an extreme heat wave false alarm (i.e., a false positive). This is not a statistical decision. It is a policy decision in which subjective values and judgments are essential. Ideally, a preferred target ratio of the number of false negatives to the number of false positives is determined before the analysis begins and is then used to help formulate the data collection and key features of the forecasting algorithm. This has worked well in other applied settings (Berk, 2012).
The conflation of corrections for sample imbalance with the estimated misclassification cost ratio is easily illustrated. Using the expressions above for the weights, the PNW forecasting exercise was repeated. Recall that the response variable was binary, and there were the same number of grid cells that experienced a heat wave as did not experience a heat wave. Using the month of June, 2021 as the target population, for a little more than 13% of the days, the 144 grid cells experienced the extreme heat wave. For the other days, the 144 grid cells did not. In the data analyzed, for 50% of the days, the 144 grid cells experienced the extreme heat wave. A difference between 13% and 50% is the kind of disjuncture addressed by Manski and Lerman (1977). Note the substantial disparity materialized even though by choosing June, 2021 and using grid cells as the study units, the imbalance was statistically manageable; the extreme heat wave was not disastrously rare.
Before weighting, the classification error rate was a little over 5%, and there were nearly the same number of false negatives as false positives. The cost ratio was about 1.0, conveying equal costs. When Manski-Lerman weighing was applied, the classification error rate rose to nearly 14%, and the cost ratio of false positives to false negatives increased from 1.0 to 9.67. False positives were now well over 9 times more costly than false negatives. In other words, the costs of heat wave false alarms were substantially up-weighted compared to the costs of true alarms not given. Many stakeholders might strongly prefer the reverse. The statistical point is that changing the prior of the response variable can dramatically change the relative costs of false negatives and positives provided by a classifier. To concentrate only on corrections for unrepresentative, response variable priors is far too narrow.
We currently favor using, when feasible, data collection and curation methods that alter the prior joint distribution of the heat wave data, as we did for the PNW and greater Phoenix area. Random forests, gradient boosting and other classifiers can respond as if the costs of false negatives and false positives are asymmetric. However, the cost ratio should be determined by stakeholders as the research is being planned. We currently are not operating under those circumstances. Without weighting, the classification cost ratios for the forecasting reported above were about 1 to 1, which may be a useful place to start the discusion. Results for other cost ratios can be computed with weighting or by altering directly the proportion of heat wave observations in the data through selective case augmentations of, or deletions from, the study dataset. The more general approach is to employ asymmetric loss functions directly, a discussion of which would take as far afield.
11 Summary and Conclusions
There are several defining features represented in our attempt to forecast extreme heat waves. To begin, within supervised learning traditions used for forecasting, the main response variable is a binary label indicating whether an extreme heat wave has occurred or not. With so many extant definitions of heat waves, we settled on the prospect of a heat dome despite its dynamic properties. There is some evidence that extreme heat waves are more than ordinary warm summer weather (Li et al., 2024).
We rely on the AIRS remote sensing data. These data do not depend on climate simulations of any kind including downscaled global climate models. AIRS data are probably noisier than simulation data and provide fewer variables with which to work; the data have promise but offer special challenges. Despite the challenges, the data in the PNW and in the Phoenix area appear to deliver instructive results.
Although our work is informed by the climate science literature, we make no explicit use of models. Our analyses rest on quasi-experimental designs, a supervised learning algorithm well suited for forecasting, and a genetic algorithm for simple feature engineering. Accurate forecasting is our primary goal. Any explanatory results are a secondary benefit.
By taking the forecasting especially seriously, forecasting uncertainty had to be taken especially seriously as well. We capitalized on recent developments in nested conformal prediction sets that can provide guarantees of valid results even in finite samples as long as the calibration data are exchangeable. We make the case for exchangeability within a particular crossover design. The split sample approach has the added benefit of working with a wide range of classifiers even if they are less than ideal by themselves.
We forecast extremely well for the Seattle area heat dome in late June of 2021 and for the Phoenix area heat dome in early July of 2023. The same predictors do well in both regions. Although the partial dependence plots are not identical in the two locations, a case can be made that they are qualitatively similar despite important differences in local climate, topography, research designs and data quality. And both sets of partial dependence plots also tend to be highly nonlinear. Still, more study sites surely are needed.
Finally, even if our findings are robust, there are many steps between hand-tailored research procedures and production quality forecasting tools that can be routinely used to deliver accurate heat dome warnings sufficiently in advance. We do not discuss such matters in this paper. That would be premature.